OpenStudy (anonymous):

The base of a parallelogram is four times its height. If the area of the parallelogram is 324 cm², whats base and height ?

OpenStudy (allank):

Hint: Area of a parallelogram = base*height Try setting the height to be an unknown. Maybe x. Can you solve now?

OpenStudy (anonymous):

I know, thanks though

OpenStudy (anonymous):

Can soomeone else help me outt?

OpenStudy (allank):

Okay, let height be x so the base will be 4x (four times the height) Area=base*height 324=4x*x=4x^2=(2x)^2 Find the square root of both sides \[\sqrt{324}=2x\] Thus 18=2x Hence x=9

OpenStudy (allank):

Thus the height = x =9, and the base = 4x = 36.